### v1

@@ 4 Bounds on $g_{3}$ coupling
where $A^{\prime}(\mu,t)=A(\mu,t)e^{\alpha\mu^{2}}$. Note that there is no subtraction term $b(t)$, otherwise the $s\leftrightarrow u$ crossing symmetry would be broken.
Moreover, if we further assume full crossing symmetry, we can derive the following expression for $a(t)$:
$a(t)=a(0)+e^{-\alpha t^{2}}t^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,0)}{\mu^{2}}\left(\frac{1}{\mu-t}+\frac{1}{\mu+t}\right)-e^{-\alpha t^{2}}t^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,t)}{\mu^{2}(\mu+t)}\,,$ (20)
where $a(0)=0$ in our case since we have $A_{\rm EFT}(0,0)=0.$ We can easily verify that in the limit $t\rightarrow 0$ the dispersion relation (19) consistently recovers (11) which was derived in the forward limit.
Equipped with a dispersion relation that is valid for a physical non-zero $t,$ we can derive constraints on additional Wilson coefficients, e.g. on the coupling $\tilde{g}_{3}$. To do so, we need to implement the partial-wave decomposition for the discontinuities in $d=4$ spacetime dimensions, i.e.
$A(s,t)=\sum_{l\,\rm{even}}n_{l}f_{l}(s)\mathcal{P}_{l}\left(1+\frac{2t}{s}\right)\,,\qquad n_{l}=16\pi(2l+1)\,,$ (21)
from which we get
where $f_{l}(s)$ is the partial-wave amplitude, $\rho_{l}(s)={\rm Im}f_{l}(s)$ is the spectral density, $\mathcal{P}_{l}(x)$ are the Legendre polynomials, and the angular momentum $l$ only takes even values because we are working with a scalar field theory. If we define
the dispersion relation (19) can be written in the compact form,
$A(s,t)=e^{-\alpha s^{2}}s^{2}\left[\frac{\mathcal{P}_{l}(1+2t/\mu)}{\mu^{2}(\mu-s)}\right]+e^{-\alpha u^{2}}u^{2}\left[\frac{\mathcal{P}_{l}(1+2t/\mu)}{\mu^{2}(\mu-u)}\right]+a(t).$ (24)
Taking derivatives with respect to $s$ and $t$ and setting $s\rightarrow 0$, $t\rightarrow 0$ afterwards, we obtain
$g_{2}=\left[\frac{1}{\mu^{3}}\right],\quad g_{3}=\left[\frac{3}{\mu^{4}}\right]-\left[\frac{2L^{2}}{\mu^{4}}\right],\quad g_{4}=\left[\frac{1}{2\mu^{5}}\right]-\left[\frac{\alpha}{2\mu^{3}}\right].$ (25)
where $L^{2}\equiv l(l+1)$. Crossing symmetry implies that the relation for the coupling $g_{4}$ in (3) can be derived in two different ways. The first way is differentiating four times with respect to $s$ and then taking $s,t\rightarrow 0$. The second way is differentiating twice with respect to $s$ and twice with respect to $t$ and then taking $s,t\rightarrow 0$. Since the two procedures must be equivalent, we get the so-called null constraint reflecting the crossing symmetry of low-energy EFT amplitude,
$\left[\frac{L^{2}(L^{2}-8)}{\mu^{5}}\right]=0\,.$ (26)
Let us make two non-trivial observations about the comparison between the method we are using to derive constraints on $g_{3}$ in non-local theories and the procedure in the case of local theories.
  • The null constraint (26) has the same form as in local theories, except for the different definition of the brackets that are adapted to the case of exponentially bounded amplitudes.
  • The relation (25) between $g_{2}$, $g_{3}$, and $g_{4}$ gets modified only through the one for $g_{4}$ (modulo brackets definition), such that the bound in the non-local case can be obtained from the local one upon the substitution $\tilde{g}_{4}^{\rm local}\rightarrow\tilde{g}_{4}+\tilde{\alpha}/2$.
The state-of-the-art technique to derive the bound on $\tilde{g}_{3}$ is based on the theory of moments and is described in [69, 70]. We briefly review the formalism in App. B and also show that it does not depend on the particular definition of the brackets. In other words, up to additional $\alpha$-dependent terms in some of the relevant relations, the constraints in the $g_{3}$-$g_{4}$ plane in our non-local model can be derived by following the same steps as in the local case [69, 70]. The result is given by the following two sets of inequalities:
@@
The most optimal bound on $g_{3}$ requires the use of the next order null constraint [69], which will certainly be modified in the non-local case. It is worth mentioning that adding higher-order null constraints would only lead to a very small improvement of the bound in its upper left corner. This does not affect any of our further conclusions and comparisons with causality bounds in the next section.
In Fig. 3a we have shown the allowed regions in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane given by the bounds (29), for different values of $\tilde{\alpha}$. The local constraint is just shifted down allowing for negative values of $g_{4}$, which are forbidden in the standard local case.
Figure 3: (a) Allowed region in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane for $\tilde{\alpha}=0,0.5,1$ from (29) and (32). (b) Allowed regions by IR causality bounds (blue region) [71] and bounds on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ in both local ($\tilde{\alpha}=0$) and non-local ($\tilde{\alpha}=1.02$) cases. For $\tilde{\alpha}<1.02$ there is an overlap between EFT constraints and the requirement of IR causality, while larger values of $\tilde{\alpha}$ are in contrast with IR causality.
5 IR causality bounds
The requirement of causality for the low-energy theory imposes a set of bounds on the Wilson coefficients of the EFT expansion (3), that must hold independently of the type of UV completion. Causality can be probed by considering a process in which a plane wave propagates on a classical background created by some local source [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. The causality condition then requires that the presence of the background should not induce superluminal propagation of the plane wave.
However, the uncertainty principle prevents us from detecting very small time advances in the signal propagation, i.e. $|\Delta t|\lesssim 1/\omega,$ where $\omega$ is the energy of the propagating wave. Therefore, unobservable causality violations are still physically allowed because, thanks to their quantum stochastic nature, they do not sum up into macroscopic time advances in signal propagation [86, 87, 71, 88]. This means that we can physically admit sufficiently short time advances that are within the uncertainty principle, i.e. time delay $\Delta t$ can also get very small negative values:
@@ 5 IR causality bounds
A comprehensive study of causality bounds in the case of shift-symmetric scalar theories based on the condition in Eq. (33) was performed in [71]. We plotted the causality constraints on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ derived in [71] together with our bounds in Fig. 3b. [^6] Authors are indebted to Marianna Carillo-Gonzales for sharing the numerical results of [71] on IR causality bounds. Let us make some important remarks on our results.
  • IR causality bounds are non-compact and are significantly weaker than local positivity constraints on the Wilson coefficients, thus leaving a place in the space of EFT parameters for causal EFTs whose UV completion can only be non-local.
  • As already noticed in the previous section, exponentially bounded amplitudes allow for negative values of $\tilde{g}_{4}$, which are not compatible with local UV completions. On the other hand, Fig. 3b shows that IR causality bounds (which are independent of the type of UV completion) allow for a very small negative $\tilde{g}_{4}$.
  • Furthermore, if the non-local scale $\alpha$ is too large, then the allowed regions derived from the non-local positivity bounds will become incompatible with the IR causality constraints. This way, we get a very interesting constraint on the UV theory imposed by IR requirements:
    $\tilde{\alpha}=\alpha\Lambda^{4}<1.02\quad\Leftrightarrow\quad M\leq(1.02)^{1/4}\Lambda\simeq\Lambda\,,$ (34)
@@
    where the second inequality is written in terms of the non-local energy scale $M\equiv 1/\alpha^{1/4}.$ This means that a causal EFT with cutoff $\Lambda$ can admit a non-local UV completion only if the scale of non-locality $M$ is of the same order or lower than the EFT cutoff.
  • On the other hand, our positivity bounds still allow for large values of the scale $\alpha$ (i.e. $\alpha\Lambda^{4}>1.02$). However, these values might not be compatible with partial-wave unitarity of the full amplitude. Although we do not have a proof, some examples supporting this statement are provided in App. A, where we notice that partial-wave unitarity is satisfied only if the EFT cutoff is close to $\alpha^{-1/4}$. Insights from IR causality might suggest this situation is generic.
@@
Finally, it is worth mentioning that non-local UV completions whose amplitudes are compatible with the properties of exponential boundedness, unitarity, analyticity and IR causality, may still show some causality violation at high energies. Our study cannot say anything about UV causality, but this is an aspect that will be investigated in future work.
6 Conclusions
@@
In this paper, we derived generalized positivity bounds on the Wilson coefficients of the EFT expansion assuming that the low-energy theory admits a non-local UV completion. In particular, we focused on a scalar theory. The central object of our discussion was a $2\rightarrow 2$ elastic scattering amplitude satisfying the properties of analyticity, crossing symmetry, and partial-wave unitarity, but not polynomial boundedness. Indeed, we considered amplitudes that are exponentially bounded in the complex energy plane, which are typical of non-local QFTs.
To compensate for the exponential growth of the amplitude we derived modified dispersion relations. While this was straightforward in the forward limit, for non-zero momentum transfer we had to require a specific but still quite generic crossing-symmetric form of the amplitude. By doing this, we successfully ensured that the integral over the arc at infinity did not diverge but vanished, thus suitable dispersion relations could be used. Consequently, we derived “non-local positivity bounds” and compared them with the standard local ones. We showed that the allowed regions in the space of EFT parameters can either overlap or not with those in the local case, depending on the value of the ratio of the non-locality scale to EFT cutoff. In particular, we found regions where the corresponding EFTs do not admit a local UV completion, but only a non-local one.
@@
We also compared the new positivity bounds with low-energy causality constraints [71]. We found that causal EFTs can admit non-local UV completions only if the scale of non-locality – defined as $M=\alpha^{-1/4}$ – is of the same order or smaller than the EFT cutoff $\Lambda$, i.e. if the following inequality is satisfied: $M<(1.02)^{1/4}\Lambda.$
@@
To support our main findings, in app. A we also presented an explicit example that serves as a proof of concept for the existence of a non-perturbative amplitude that satisfies all the standard desired properties (i.e. partial-wave unitarity, causality, crossing symmetry) except for the polynomial boundedness.
@@
It is also worth to mention that non-local (or non-localizable) theories can be relevant for the construction of a UV completion of general relativity in the form of four-dimensional QFTs, such as infinite derivative gravity [91], or asymptotically safe gravity formulated in terms of form factors in the quantum effective action [92, 93]. Indeed, it would also be interesting to compare our positivity bounds with constraints derived in the context of asymptotically safe theories [94, 95, 96, 97]. Our work reinforce the idea that these quantum-gravity approaches could be consistent with the full set of unitarity constraints and with macroscopic causality. Exploring and classifying different types of viable exponentially bounded amplitudes could be a very interesting direction for future studies. This has the potential to shed new light on the form of the quantum effective action in UV-complete gravitational theories, as well as on the question of the uniqueness of string theory as the only consistent UV completion of general relativity.
Acknowledgements
@@
Authors are indebted to Ivano Basile, Gia Dvali, Alessia Platania, Xinan Zhou, and Shuang-Yong Zhou for several enlightening conversations. A. T. thanks Alexander Zhiboedov for his hospitality at CERN and numerous deep discussions on the reconstruction of the unitary and causal amplitudes. A. T. thanks GuangZhuo Peng and YoungJun Xu for useful comments and feedback. The authors are grateful to Marianna Carillo-Gonzales for sharing her numerical results on IR causality bound. L. B. and A. T. thank The Royal Society for financial support through the International Exchanges Scheme 2022. L. B. acknowledges financial support from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Actions (grant agreement ID: 101106345-NLQG). The work of A. T. was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 1234710.
Appendix A Example of unitary and causal amplitude violating Froissart-Martin bound
Scattering amplitudes that are exponentially bounded in the complex energy plane can be obtained, for example, in theories containing infinite derivative operators [98]. This type of non-polynomial behavior is usually associated to the lack of locality of the underlying field theory. It is still under debate whether these theories are consistent with the other standard fundamental requirements, such as causality. Here we construct working examples of UV complete amplitudes that are exponentially bounded but still satisfy partial-wave unitarity, crossing symmetry and causality.
The simplest form of such an amplitude could be
$A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}\,.$ (35)
However, it can be shown that this expression does not satisfy partial-wave unitarity for large $s$ due to the fact that the corresponding partial waves heavily oscillate and can give rise to negative imaginary contributions. In other words, if we compute
${\rm Disc}A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}$ (36)
and expand in partial waves, it can be shown that the imaginary part of the partial waves can be negative for large values of $s.$
However, we can try to suitably modify the amplitude and check whether we can find expressions that are compatible with unitarity. For example, it can be shown that the following discontinuity,
${\rm Disc}_{s}A(s,t,u)\propto-\left(e^{-\alpha s^{2}}\log{s}+e^{-\alpha t^{2}}\log{t}+e^{-\alpha u^{2}}\log{u}\right)\,,$ (37)
has positive-definite $s$-discontinuities of the partial waves. A possible form of the amplitude whose discontinuity is of the type in eq. (37) but is defined in the physical region (i.e. $t\leq 0$) is given by
$A(s,t,u)=$ $C\left[e^{-\alpha s^{2}}\log{s}\log{(-s)}+e^{-\alpha t^{2}}\log{t}\log{(-t)}+e^{-\alpha u^{2}}\log{u}\log{(-u)}\right]$ (38)
$=$ $C\left[e^{-\alpha s^{2}}\log{(s)}\,(\log{s}-i\pi)+e^{-\alpha t^{2}}\log{(-t)}\,(\log{(-t)}-i\pi)+\right.$
$\qquad\qquad\qquad\left.+e^{-\alpha u^{2}}\,\log{(-u)}(\log{(-u)}-i\pi)\right]\,.$
Here the branch cut of the logarithm is specified by the condition $\log{s}-\log{(-s)}=i\pi$. This amplitude can be shown to satisfy partial-wave unitary for reasonable values of the pre-factor $C.$ The behavior of the real and imaginary parts of some partial waves are plotted in Fig. 4a for $C=1$.
A weak point the amplitude (38) is that some of the partial waves blow up in the $s\rightarrow 0$ limit (see Fig. 4a). We can still find a working example with all imaginary parts of the partial waves that are positive and finite. This is given by the amplitude
$A(s,t,u)=C\left(s^{\gamma i}e^{-\alpha s^{2}}+t^{\gamma i}e^{-\alpha t^{2}}+u^{\gamma i}e^{-\alpha u^{2}}\right)\,.$ (39)
If $\gamma$ is small, the power can be expanded in terms of logarithms, so it will behave similarly to (38). In what follows we show that for this amplitude unitarity and causality requirements are satisfied for all real positive $s$. Therefore, this can serve as a proof of concept for the existence of a consistent exponentially bounded amplitude that is compatible with the crossing symmetric structure in (18), which was used in this paper to derive positivity bounds.
@@ Appendix A Example of unitary and causal amplitude violating Froissart-Martin bound
$f_{l}\left(s\right)=\dfrac{1}{16\pi}\int^{1}_{-1}{\rm d}xP_{l}\left(x\right)A\left(s,-\dfrac{s}{2}\left(1-x\right),-\dfrac{s}{2}\left(1+x\right)\right).$ (41)
Figure 4: (a) Real (solid lines) and imaginary (dashed lines) parts of the partial-wave amplitudes as functions of $s$ for the amplitude (38). (b) Real (solid lines) and imaginary (dashed lines) parts of the partial-wave amplitudes as functions of $s$ for the amplitude (39) with $\gamma=0.1$. We set $C=1$ in both figures.
For the amplitude (39), we can write an analytic series expression of the partial waves for $l>0$ by expanding the exponent and integrating term by term, i.e.
$f_{l}(s)=\dfrac{1}{4\pi}\sum^{\infty}_{n=0}\frac{\left(-\alpha\right)^{n}\left(-s\right)^{2ng+i\gamma}}{\Gamma(n+1)}\dfrac{\left(\Gamma\left(i\gamma+2ng+1\right)\right)^{2}}{\Gamma\left(i\gamma+2ng+l+2\right)\Gamma\left(i\gamma+2ng-l+1\right)}.$ (42)
From this expression it is evident that the partial waves have the same type of essential singularity at infinity of the full amplitude, that is, they are not polynomially bounded in a complex energy plane. However, the unitarity condition is required only for physical momenta $s\geq 0$, which means that their growth in the complex energy plane is not necessarily in contradiction with partial-wave unitarity.
The sum in the Eq. (42) can be expressed in terms of rather complicated hypergeometric functions, but we noticed that it was more efficient to check unitarity by computing partial waves numerically from (41). We found that the amplitude (39) satisfies the positivity requirement ${\rm Im}f_{l}(s)>0$ if $\gamma\leq 0.1$. As all partial waves are bounded, the full non-linear unitarity condition can be always achieved by suitably adjusting the overall constant factor. In Fig. 4b we plotted the numerical solutions for $\gamma=0.1.$
@@
$\Delta T=2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}\,\delta_{l}(s)\approx 2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}f_{l}(s)>0\,,\qquad f_{l}(s)=-i(e^{i\delta_{l}(s)}-1).$ (43)
Here $\delta_{l}$ is a scattering phase corresponding to the partial wave amplitude, and the approximation used means that $f_{l}(s)$ is small.
Due to oscillations in the real part of the partial waves for small values of $s$ (see Fig. 4b), this strict causality condition cannot be satisfied. However, the condition of macroscopic causality that still allows for unresolvable time advances (33) can instead be respected. This condition is more justified from a quantum mechanical point of view and is given by

### v2

@@
where $A^{\prime}(\mu,t)=A(\mu,t)e^{\alpha\mu^{2}}$. The expression for $a(t)$ can be derived by imposing full crossing symmetry and we have $a(0)=0$. Note that, there is no subtraction term $s\,b(t)$, otherwise the $s\leftrightarrow u$ crossing symmetry would be broken.
Figure 3: (a) Allowed region in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane for $\tilde{\alpha}=0,0.5,1$ from (16) and (19). (b) Allowed regions by IR causality (blue region) [70] and positivity bounds on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ in both local ($\tilde{\alpha}=0$) and non-local ($\tilde{\alpha}=1.02$) cases. For $\tilde{\alpha}<1.02$ there is an overlap between constraints from positivity bounds and IR causality, while larger values of $\tilde{\alpha}$ are in contrast with IR causality.
The state-of-the-art technique to derive the bound on $g_{3}$ is based on the theory of moments and is described in [69, 71]. We briefly review the formalism in the Supplemental Material and also show that it does not depend on the particular definition of the brackets. In other words, up to additional $\alpha$-dependent terms in some of the relevant relations, the constraints in the $g_{3}$-$g_{4}$ plane in our non-local model can be derived by following the same steps as in the local case [69, 71]. The result is given by the following two sets of inequalities:
@@
In Fig. 3a we have shown the allowed regions in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane given by the bounds (16) and (19), for different values of $\tilde{\alpha}$. The local constraint is just shifted down allowing for negative values of $g_{4}$, which are forbidden in the standard local case.
IR causality bounds.— The requirement of causality for the low-energy theory imposes a set of bounds on the Wilson coefficients of the EFT expansion (3), that must hold independently of the type of UV completion. Causality can be probed by considering a process in which a plane wave propagates on a classical background created by some local source [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 70, 83, 84], and its validity demands the absence of superluminal propagation. However, the uncertainty principle prevents us from detecting very small time advances in the signal propagation, i.e. $|\Delta T|\lesssim 1/\omega,$ where $\omega$ is the energy of the propagating wave. Therefore, unobservable causality violations could still be physically allowed because, thanks to their quantum stochastic nature, they do not sum up into macroscopic time advances in the signal propagation [85, 80, 70, 86]. This means that we can physically admit sufficiently short time advances that are within the uncertainty principle, i.e. time delays $\Delta T$ can also get very small negative values:
@@
A comprehensive study of causality bounds in the case of shift-symmetric scalar theories based on the condition in Eq. (20) was performed in [70]. We plotted the causality constraints on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ derived in [70] together with our bounds in Fig. 3b. We find the following results. (i) IR causality bounds are non-compact and significantly weaker than local positivity constraints on the Wilson coefficients, thus leaving a place in the space of EFT parameters for causal EFTs whose UV completions can only be non-local. (ii) Exponentially bounded amplitudes allow for negative values of $g_{4}$ that are compatible with IR causality and a non-local UV completion. (iii) A causal EFT with cutoff $\Lambda$ can admit a non-local UV completion only if the scale of non-locality is of the same order or larger than the EFT cutoff, i.e.
$\tilde{\alpha}=\alpha\Lambda^{4}<1.02\quad\Leftrightarrow\quad M\geq(1.02)^{-1/4}\Lambda\simeq\Lambda\,,$ (21)
@@
where the second inequality is written in terms of the non-local energy scale $M\equiv 1/\alpha^{1/4}.$
@@
It is worth to mention that UV completions whose amplitudes are compatible with exponential boundedness, unitarity, analyticity and IR causality, may in principle still show some causality violation at high energies. Our study cannot say anything about UV causality, but this is an aspect that will be investigated in future work.
However, we have managed to construct exponentially bounded, unitary amplitudes that satisfy UV causality, in particular the condition of asymptotic causality. An example is given by $A(s,t,u)=C+\left[F(s)+F(t)+F(u)\right]$, where
$F(z)=-\sqrt{\alpha}z\,_{2}F_{2}\left(\frac{1}{2},\frac{1}{2};\frac{3}{2},\frac{3}{2};-\alpha z^{2}\right)+\frac{\sqrt{\pi}}{2}\,\text{erf}\left(\sqrt{\alpha}z\right)\left[\log\left(\sqrt{\alpha}z\right)+i\pi\right],$ (22)
and $C$ is some constant. In the Supplemental Material we explicitly show that partial-wave unitarity and asymptotic causality are satisfied by this amplitude.
@@
Conclusions.— In this Letter, we derived generalized positivity bounds on the Wilson coefficients of the EFT expansion, assuming that the low-energy theory admits a non-local UV completion according to which the amplitudes are exponentially bounded in the complex $s$ and $t$ planes.
@@
We showed that the allowed regions in the space of EFT parameters can either overlap or not with those in the local case, depending on the value of the ratio of the non-locality energy scale to EFT cutoff. In particular, we found regions where the corresponding EFTs do not admit a local UV completion, but only a non-local one. We also compared the new positivity bounds with low-energy causality constraints [70] and found that causal EFTs can admit non-local UV completions only if the scale of non-locality – defined as $M=\alpha^{-1/4}$ – is of the same order or larger than the EFT cutoff $\Lambda$.
@@
To support our main findings, we provided an explicit example that serves as a proof-of-concept for the existence of amplitudes that satisfy all the standard desired properties (i.e. partial-wave unitarity, causality, crossing symmetry) except for the polynomial boundedness.
@@
Understanding non-localizable amplitudes can be relevant for describing field-theoretic aspects of black-hole formation via high-energy scattering [89, 68], and for formulating QFT-based UV completions of general relativity, such as infinite-derivative gravity [90] or asymptotically safe gravity [91]. Indeed, it would also be interesting to compare our positivity bounds with constraints derived in the context of asymptotically safe theories [92, 93, 94, 95]. Exploring and classifying different types of viable exponentially bounded amplitudes could be an interesting direction for future studies. This has the potential to shed new light on the form of the quantum effective action in UV-complete gravitational theories, as well as on the question of the uniqueness of string theory as the only consistent UV completion of general relativity.
@@
Acknowledgements.— The authors are indebted to Ivano Basile, Gia Dvali, Alessia Platania, Xinan Zhou, and Shuang-Yong Zhou for several enlightening conversations, and are grateful to Marianna Carillo-Gonzales for sharing her numerical results on IR causality bounds from Ref. [70]. A. T. thanks Alexander Zhiboedov for his hospitality at CERN and numerous deep discussions on the reconstruction of unitary and causal amplitudes. A. T. thanks GuangZhuo Peng and YoungJun Xu for useful comments and feedback. L. B. and A. T. thank The Royal Society for financial support through the International Exchanges Scheme 2022. L. B. acknowledges financial support from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Actions (grant agreement ID: 101106345-NLQG). The work of A. T. was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 1234710.
Supplemental material
Unitary and causal amplitudes violating Froissart-Martin bound
Here we construct working examples of UV complete amplitudes that are exponentially bounded but still satisfy partial-wave unitarity, crossing symmetry and (asymptotic) causality.
@@
$f_{l}\left(s\right)=\dfrac{1}{16\pi}\int^{1}_{-1}{\rm d}x\mathcal{P}_{l}\left(x\right)A\left(s,-\dfrac{s}{2}\left(1-x\right),-\dfrac{s}{2}\left(1+x\right)\right).$ (24)
In this work, we mainly focused on crossing-symmetric amplitudes of the form
$A(s,t,u)=F(s)+F(t)+F(u)\,.$ (25)
Assuming that $F(z)$ can be expanded in series around zero, we can write
$F(z)=z^{\beta}\sum_{n=0}^{\infty}a_{n}z^{2gn},$ (26)
where $\beta$ can also be non-integer. Consequently, we can write an analytic expression of the partial waves for $l>0$ by integrating the latter series term by term, i.e.
$f_{l}(s)=\dfrac{1}{4\pi}\sum^{\infty}_{n=0}a_{n}\left(-s\right)^{2ng+\beta}\dfrac{\left[\Gamma\left(\beta+2ng+1\right)\right]^{2}}{\Gamma\left(\beta+2ng+l+2\right)\Gamma\left(\beta+2ng-l+1\right)}.$ (27)
For exponentially bounded amplitudes with $F(s)\sim e^{-\alpha s^{2g}}$ we have $a_{n}=(-\alpha)^{n}/n!$. In this case, from the expression (27) it is evident that the partial waves have the same type of essential singularity at infinity ($f_{l}(s)\sim e^{-\alpha s^{2g}}$ at large $s$) as that of the full amplitude, that is, they are not polynomially bounded in the complex $s$-plane. However, the unitarity condition is required only for physical momenta $s\geq 0$, which means that their growth in the complex $s$-plane is not necessarily in contradiction with partial-wave unitarity.
The sum in the Eq. (27) can be expressed in terms of rather complicated hypergeometric functions, but we noticed that it was more efficient to check unitarity by computing partial waves numerically from (24). As all partial waves are bounded, the full non-linear unitarity condition can be always achieved by suitably adjusting the overall constant factor.
The simplest form of an exponentiall bounded amplitude could be
$A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}\,.$ (28)
However, it can be shown that this expression does not satisfy partial-wave unitarity for large $s$ due to the fact that the corresponding partial waves heavily oscillate and can give rise to negative imaginary contributions. In other words, if we assume
${\rm Disc}_{s}A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}$ (29)
and expand in partial waves, it can be shown that the imaginary parts of the partial waves can be negative for large values of $s$.
However, we can try to modify the amplitude suitably and check whether we can find expressions that are compatible with unitarity. For example, the amplitude
${\rm Disc}_{s}A\propto\left(e^{-\alpha s^{2}}\log{s}+e^{-\alpha t^{2}}\log{t}+e^{-\alpha u^{2}}\log{u}\right),$ (30)
has positive-definite $s$-discontinuities of the partial waves. A possible form of the amplitude with this type of discontinuity, but defined in the physical region (i.e. $t\leq 0$), is given by
$A(s,t,u)=C\left[e^{-\alpha s^{2}}\log{s}\log{(-s)}+e^{-\alpha t^{2}}\log{t}\log{(-t)}+e^{-\alpha u^{2}}\log{u}\log{(-u)}\right].$ (31)
Here the branch cut of the logarithm is specified by the condition $\log{s}-\log{(-s)}=i\pi$. This amplitude can be shown to satisfy partial-wave unitary for reasonable values of the pre-factor $C$. The behavior of the real and imaginary parts of the first partial waves are plotted in Fig. 4a for $C=1$ and $\alpha=1$.
Figure 4: Real (solid lines) and imaginary (dashed lines) parts of the partial-wave amplitudes as functions of $s$ for the amplitudes (31) (left panel (a)) and (32) (right panel (b)). We set $\alpha=1$ and $\gamma=1$.
A weak point of the amplitude (31) is that some of the partial waves blow up in the $s\rightarrow 0$ limit (see Fig. 4a). Is there a possibility to write an amplitude with finite partial waves in the IR limit? Apparently, it is relatively easy to construct a unitary amplitude that is exponentially bounded. For example, the choice
$F(z)=-i\,\Gamma(\gamma,\alpha z^{2})\log{\sqrt{\alpha}z}\,,\qquad\gamma>0,$ (32)
where $\Gamma(\gamma,z)$ is the incomplete gamma-function, provides an amplitude whose partial waves have positive and bounded imaginary parts, as shown in Fig.4b.
@@
$\sqrt{s}\Delta T=2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}\,\delta_{l}(s)\approx 2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}f_{l}(s)>0\,,$ (33)
Here $\delta_{l}$ is a scattering phase corresponding to the partial-wave amplitude written as
$f_{l}(s)=-i(e^{i\delta_{l}(s)}-1)\,,$ (34)
and the approximation used means that $f_{l}(s)$ is small.
Figure 5: (a) Real (solid lines) and imaginary (dashed lines) parts of the partial wave $l=0$ for the amplitude (35) as a function of $s$. (b) Imaginary parts of the partial waves $l=2,4,6,8$ for the amplitude (35) as a function of $s$. We set $C=5i/2$. All partial waves have positive imaginary parts. The real part of $f_{0}(s)$ is a monotonically growing function which guarantees that the asymptotic causality condition (33) is satisfied.
Oscillations in the real part of the partial waves can violate the condition of asymptotic causality, indeed this happens for the amplitude (32). However, we can still find amplitudes that satisfy both partial-wave unitarity and causality. An explicit example is by
$A(s,t,u)=C+\left[F(s)+F(t)+F(u)\right]\,,$ (35)
where
$F(z)=$ $\int_{0}^{z}{\rm d}y\log{\left(-\sqrt{\alpha}y\right)}e^{-\alpha y^{2}}$ (36)
$=$ $-\sqrt{\alpha}z\,_{2}F_{2}\left(\frac{1}{2},\frac{1}{2};\frac{3}{2},\frac{3}{2};-\alpha z^{2}\right)+\frac{1}{2}\sqrt{\pi}\,\text{erf}\left(\sqrt{\alpha}z\right)\left[\log\left(\sqrt{\alpha}z\right)+i\pi\right],$
and $C$ can be an arbitrary constant. It is important to mention that the presence of a non-zero constant $C$ does not affect any of the dispersion relations used in this paper since the relevant results were obtained by taking derivatives with respect to $s$ and $t$.
Partial-wave amplitudes corresponding to (35) are plotted in Fig. 5 from which we clearly see that the real parts are monotonically growing functions of $s$ (i.e. there is no oscillation). Note that the only partial wave that has non-zero real part is $l=0.$ The constant term $C$ in (35) is therefore required for the amplitude so that the $l=0$ partial wave satisfies the asymptotic causality.
This last example provides a proof-of-concept for the existence of amplitudes that are non-localizable (i.e. exponentially bounded) but, at the same time, satisfy partial-wave unitarity and asymptotic causality (i.e. absence of time advances).
Let us remark that the requirements of unitarity and analyticity are not enough to guarantee the validity of asymptotic causality. In fact, the weaker condition of macroscopic causality, which allows for unresolvable time advances could be more justified from a quantum mechanical point of view. The latter is given by